OpenStudy (anonymous):

2x+10/ (x+5)^2 + x/x+5??

OpenStudy (austinl):

What exactly do you need done with this? Simplified?

OpenStudy (anonymous):

Add and simplify austinL

OpenStudy (austinl):

Well, you are adding two unlike fractions so you need to get a like denominator first. Can you do that?

OpenStudy (anonymous):

Would it be x+2/x+5?

OpenStudy (austinl):

For a final answer?

OpenStudy (anonymous):

Ya

OpenStudy (austinl):

That is not what I got for a final answer.

OpenStudy (austinl):

What you need to do is get a common denominator by multiplying the left side of the equation, top and bottom, by the denominator on the right. And vice versa for the right.

OpenStudy (austinl):

You do that?

OpenStudy (anonymous):

\[2x+10/ (x+5)^2 + x/x+5 \]has a lowest possible common denominator of \[(x+5)^2\] To get all of them to have that denominator we have to multiply each term sufficiently by "1"\[(2x)*\frac{ (x+5)^2 }{(x+5)^2} \]gives us the correct form for our first part without changing its value. \[\frac{ 10 }{ (x+5)^2 }\]is already in the form we need and \[\frac{ (x) }{ (x+5) }* \frac{ (x+5) }{ (x+5) }\] should do the trick. now we have the correct form for each piece and can thus add them together. Note that we have not added or removed any value, we have simply multiplied by "1" and hence have changed the form. This results in:\[\frac{ [(2x)(x+5)^2 + 10 +(x)(x+5)]}{ (x+5)^2 }\]The answer is easy to find, you simply now need to multiply the top out and simplify where necessary :)

OpenStudy (austinl):

Once you do that, you can check your answer with me.